OpenStudy (anonymous):

solve show work. -3(7x-6)-x=26-2(11x+4) help pleez i wrote it wrong at first

OpenStudy (anonymous):

-21x+18-x=26-22x-8 (Expand) 18=18 (Move all x's onto left side) True